Let the demand for coal be Q = 100 - 10P + 20Y, where P is the price of coal and Y is income. Let the supply of coal be Q = 5P + 1G, where G is the price of natural gas from coalbed methane. If G increases, what happens to the equilibrium price and quantity of coal?

The price and quantity both increase.

The price decreases but the quantity increases

The price increases but the quantity decreases.

The price and quantity both decrease.
Answer #1

L). In equilibrium Demand = Supply 100-10P+ 20y = SP + 1 h 207 + 100 = ISP tu o = 15 d P ti du dl S du 15 decuase - 15 Price